OpenStudy (anonymous):

Use a quadratic equation to find two real numbers that satisfies the situation. The sum of the two numbers is 12, and their product is -28.

OpenStudy (asevilla5):

Well, two numbers x1 and x2 which have a sum = S and a product = P are THE solutions of equation : x^2 - Sx + P = 0 x^2 - 12x - 28 = 0 the quadratic formula gives: ∆ = (-12)^2 - 4*1*(-28) = 144+112 = 256 = 16^2 the solutions are therefore: x1 and x2 = (12 +/- 16)/2 = 14 and - 2